Nick has 3 times the money Stacey has and Lindsey has 4 times thes money that stacey has. If Rick's $120.00 is added to the group the average amt. of money the four have is $150. How much money does each person have?
Answer by rfer(16322)

n=3x=$180
s=x=$60
l=4x=$240
r=$120
-----------
(3x+x+4x+120)/4=150
8x+120/4=150
8x+120=150*4
8x+120=600
8x=480
x=60